OpenStudy (anonymous):

Determine the open intervals in which the function f(x)=3x^2-x^3 is concave upwards or downwards

OpenStudy (anonymous):

So I got the no x=1 for the 2nd derivative. on the left side, all no. are posetive. on the right. all are negative. now what does that mean?

OpenStudy (unklerhaukus):

f(x) is concave upwards , when f''(x)<0 f(x) is concave downwards , when f''(x)>0

OpenStudy (anonymous):

oh yeah! I forgot about that! Thank you!

OpenStudy (unklerhaukus):

\[f(x)=3x^2-x^3\]\[f'(x)=6x-3x^2\]\[f''(x)=6-6x\]

OpenStudy (anonymous):

Yup, got that. thank you